what is the efficiency of pump if 1500 watt and work output 12000 J in 10 seconds
1
Expert's answer
2019-09-12T10:40:56-0400
"\rm efficiency=\frac{output\: work}{input\: work}"

"\eta=\frac{W_{output}}{W_{input}}=\frac{12000\:\rm J}{1500\:\rm W\times 10\:\rm s}=0.8=80\%"